# MMASD: A Multimodal Dataset for Autism Intervention Analysis

### v1

@@ 6. Conclusion
Autism research has been greatly facilitated by machine learning techniques, which offer cost-effective and accurate ways to analyze various aspects of children’s behavior and development. However, the lack of open-access datasets has posed challenges to conducting fair comparisons and promoting sound research practices in the field of autism research. To address this issue, we have proposed MMASD, a publicly accessible multimodal dataset that features diverse scenes collected from therapy interventions for children with autism. Our dataset includes multimodal features such as 2D & 3D skeleton data, optical flow, demographic data, and ADOS rating, offering a confidential data-sharing approach that can maintain critical motion information. MMASD distinguishes itself from existing works by utilizing privacy-preserving multimodal features to provide comprehensive representations of full-body movements across a range of diverse themes. Moreover, each scene in our dataset depicts the same activity performed by a child and one or more therapists, providing a valuable template for comparing typically developing individuals with children with ASD.
There are several directions for future work that are worth exploring. Firstly, further research can be conducted to develop and compare machine learning models on the MMASD dataset for various tasks, such as action quality assessment (Li et al., 2021a), interpersonal synchrony estimation (Li et al., 2022a, b), and cognitive status tracking. This can help establish benchmark performance and identify state-of-the-art methods for analyzing the full-body movements of children during therapeutic interventions. In addition, new approaches can be investigated to overcome pose detection failures in MMASD. For example, by introducing pose uncertainty or attention mechanism to assign higher weights to more reliable body joints. Furthermore, the MMASD dataset can be expanded in several aspects. To elaborate, further intervention scenarios could be included along with more features such as mutual gaze (Guo et al., 2023) and additional annotations like a synchrony score. Finally, efforts can be made to dataset augmentation via existing benchmarks not limited to the autism domain by matching samples with similar motion features (Pandey et al., 2020), which can significantly expand the scale of the autism dataset.

### v3

@@ 6. Conclusion
Autism research has been greatly facilitated by machine learning techniques, which offer cost-effective, non-invasive, and accurate ways to analyze various aspects of children’s behavior and development. However, the lack of open-access datasets has posed challenges to conducting fair comparisons and promoting sound research practices in the field of autism research. In this paper, we have proposed the MMASD, a privacy-preserving publicly accessible multimodal ASD children dataset $(N=32)$. The dataset features diverse, hand-annotated clips from over $100hrs$ of raw videos from the play therapy interventions. Our dataset includes multimodal features such as 2D & 3D skeleton data, optical flow, demographic data, and clinical rating, offering a confidential data-sharing approach that can maintain critical full-body motion information. Moreover, each scene in our dataset depicts the same activity performed by a child and one or more therapists, providing a valuable template for comparing typically developing individuals with children with ASD. The open-access MMASD dataset distinguishes itself from existing works by utilizing privacy-preserving multimodal features to provide comprehensive representations of full-body movements across diverse therapeutic social activities.
Acknowledgements.
We would like to express our sincere gratitude for our research team, therapists and students who conducted the therapy sessions (Sudha Srinivasan, Maninderjit Kaur, and Isabel Park). We also thank therapy participants and their caregivers for contributing to the MMASD dataset collection. This work is partly supported by our sponsors, Amazon Research Awards Program, University of Delaware AI Center of Excellence, the National Institute of General Medical Sciences (NIGMS, P20 GM103446E), the National Institutes of Mental Health (NIMH, 5R21MH089441-02, 4R33MH089441-03), and Autism Speaks (Grant #8137). Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of the sponsors.